This solve by substitution method is really confusing me I really need help with this please
2x+2y=2
-8x+y=28

Answer by rfer(16322)

y=8x+28
2x+2y=2
2x+2(8x+28)=2
2x+16x+56=2
18x=-54
x=-54/18
x=-3
-----------------
2*-3+2y=2
-6+2y=2
2y=8
y=8/2
y=4
---------------
-8(-3)+4=28
24+4=28
I hope that helps.
Bob